OpenStudy (anonymous):

Let u, v and w be vectors in 3d space. prove: if u and v are both orthogonal to w, then the vector u + v is also orthogonal to w.

OpenStudy (zarkon):

\[u\cdot w=0\] \[v\cdot w=0\] \[(u+v)\cdot w=u\cdot w+v\cdot w=0+0=0\]